OpenStudy (anonymous):

What is the result of adding the system of equations? 2x + y = 4 3x - y = 6

OpenStudy (abdullahm):

Hi, add the system of equations. 2x + y = 4 3x - y = 6 ___________ 2x + 3x + y + (-y) = 4 + 6 ^ simplify that